|8x+2|+5=9
Ansswer choices:
A.) {-1/4, 3/4}
B.) {1, -3}
C.) {1/4, -3/4}
D.) no solution
* I know that 1/4 is part of the answer but I'm not sure if -3/4 is.

Answer by scott8148(6628)

│8x+2│ = 4

8x+2 = ±4

8x = -6 ___ x = -3/4

8x = 2 ___ x = 1/4